# Vulnerability of Text-Matching in ML/AI Conference Reviewer Assignments to Collusions

### v1

@@ 4 Attack Procedure
For any colluding paper-reviewer pair $(p,r)$, the goal of the attack is to have $r$ be assigned to $p$ for peer review at a conference. We begin by detailing the reviewer’s adversarial action in Section 4.1; then, we detail the author’s adversarial action in Section 4.2. We provide an overview of the procedure in Figure 1.
@@
(b) The colluding author adversarially modifies their original abstract to be more similar to the paper in the colluding reviewer’s curated archive.
Figure 1: An illustrated example of the attack procedure.
@@ 4.1 Adversarial Archive Curation by the Colluding Reviewer
Peer review assignment systems often allow reviewers to curate their own archives, which are used to compute similarity scores for assigning papers. Some conferences even encourage this practice, starting with an empty archive that reviewers populate with any papers they choose. For instance, platforms like OpenReview and the Toronto Paper Matching System (TPMS) [CZ13b] allow reviewers to manually add arbitrary publications by specifying titles and abstracts, or remove articles from their current profiles. While the ability to manually add arbitrary publications can amplify the effectiveness of attacks, our analysis focuses on the more common case where reviewers select papers from their existing profiles.
A colluding reviewer can exploit this process by constructing an adversarial archive, $\mathcal{Q}_{r}^{\text{adv}}$, that includes only papers from their archive $\mathcal{Q}_{r}$ which are highly similar to the target paper $p$. With max pooling, a commonly used aggregation strategy (Section 2.2), the reviewer only needs to ensure that their archive retains the paper most similar to $p$. Since max pooling considers only the highest similarity score, the presence of one highly similar paper is sufficient to bias the assignment, making this method particularly vulnerable.
@@
$\mathcal{Q}_{r}^{\text{adv}}=\left\{\text{a random paper in }\operatorname*{argmax}_{q\in\mathcal{Q}_{r}}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}\right\},$ (1)
@@
For simplicity, the majority of our analyses focuses on the setting where the colluding reviewer has a single paper in their archive. Under max pooling (with arbitrary length archives), the similarity score obtained by the colluding reviewer for the target paper is greater than or equal to the similarity score obtained under our setting of having a single paper in the archive. Thus, while all our analyses of a single paper aren’t identical to max pooling, they give an indication of its potential under both max pooling and under reviewer-driven sub-selection under average pooling. We rigorously compare the two aggregation strategies in Section 5.6 where we find that max pooling is indeed more vulnerable.
@@ 4.2 Adversarial Abstract Modifications by the Colluding Author
Recall that only titles and abstracts are used to compute similarity scores between papers, and reviewers can curate their archives adversarially. In this section, we will expand notation for paper $p$ to have corresponding $t_{p}$ and $a_{p}$ as we will manipulate those, and we will explicitly denote the reviewer archive. When the archive has only one paper, max and average pooling are the same and we drop the subscript for brevity in notation. Hence, we use $s((t_{p},a_{p}),\mathcal{Q}_{r})$ to denote paper-reviewer similarity below.
@@
To have a paper $p$ with the title denoted by $t_{p}$ and the abstract denoted by $a_{p}$ assigned to the colluding reviewer $r$ for peer review, the key idea behind the authors’ steps is to construct for the paper $p$ an adversarial abstract $a^{\text{adv}}_{p}$ to increase the similarity $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$ from $s((t_{p},a_{p}),\mathcal{Q}_{r}^{\text{adv}})$ while maintaining consistency and coherence to avoid arousing suspicion.
We discover two types of abstract modifications that effectively increase the similarity and formalize them into two simple operations. The first is IncludeThemes which involves adding background or filler sentences related to the main ideas of papers in $\mathcal{Q}_{r}^{\text{adv}}$. The second is InsertKeywords which inserts “keywords” that target SPECTER similarity to increase $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$ even if the keywords do not necessarily seem important to humans. This operation is inspired by works in adversarial robustness that show how small unexpected and unintelligible changes can break machine learning classifiers [ERLD17, JJZS20, LMG+20].
For each operation, we provide a description of the operation and an example modified abstract. Both operations employ Large Language Models (LLMs) and have two modes in which they can operate — human-in-the-loop and fully automatic. The human-in-the-loop mode is the way we expect colluding authors would modify their abstracts in reality, as human oversight can ensure there are no suspicious artifacts in the adversarial abstract edited by LLMs. For several experiments in this paper, we use the fully automatic mode for scalability, but we also simulate and evaluate human-in-the-loop attacks at a feasible scale. The LLM model we use in this paper is the OpenAI gpt-4-0125-preview model.
@@
The goal of IncludeThemes is to modify the abstract $a_{p}$ in order to increase the SPECTER similarity between paper $p$ and reviewer $r$ by adding background or filler sentences related to the central themes of papers in $\mathcal{Q}_{r}^{\text{adv}}$. We note that the resulting modified abstract $a^{\text{adv}}_{p}$ may become inconsistent with its paper $p$ if the main ideas in the abstract have been changed entirely. Our key observation is that SPECTER similarities between papers and reviewers increase when abstracts share overlapping themes, even when those themes are central to papers in $\mathcal{Q}_{r}^{\text{adv}}$ but only referenced as background information in $a^{\text{adv}}_{p}$. This means that IncludeThemes can produce $a^{\text{adv}}_{p}{}$ by including themes from $\mathcal{Q}_{r}^{\text{adv}}$ to increase $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$ in a way that does not break the coherence and consistency constraints. An example of IncludeThemes can be found in Figure 2.
@@
Figure 2: An example of IncludeThemes modifying the abstract. Goal navigation is an important theme in the adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$ of the reviewer $r$. The modification (highlighted sentence) adds robot goal navigation as a motivating example for further improvements to multimodal models. The resulting abstract $a^{\text{adv}}_{p}$ remains coherent and consistent with the paper’s focus on text and image embedding alignment, but has increased similarity to the target adversarial archive $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$. The IncludeThemes changes shown in this example, alongside adversarial reviewer archive curation, increase the reviewer’s similarity to the paper from being 101st most-similar to become the most similar amongst all reviewers at the NeurIPS 2023 conference.
The implementation of the human-in-the-loop and fully automatic modes differ in whether human supervision is used to ensure the coherence and consistency of the modified abstract $a^{\text{adv}}_{p}$. In the human-in-the-loop mode, modified abstracts $a^{\text{adv}}_{p}$ are created by a human (potentially with the help of a LLM), and the human can make incremental edits to $a^{\text{adv}}_{p}$ to ensure coherence and consistency. On the other hand, the fully automatic mode only uses a LLM to generate the $a^{\text{adv}}_{p}$ with a prompt that emphasizes coherence and consistency, and it does not allow further edits to be made to $a^{\text{adv}}_{p}$ once they are generated.
In addition, the implementation for both human-in-the-loop and fully automatic modes involves generating $\bm{N}$ different versions of modified abstracts. Since generating and modifying abstracts is stochastic (due to the stochasticity of LLM outputs and manual edits), we keep only the attempt that is the most similar to the colluding reviewer $r$. We tune and report the choice of $\bm{N}$ for our experiments in Section 5.1, and the formal algorithm for the IncludeThemes operation are detailed in Appendix A.1.
@@
In order to further raise the similarity, InsertKeywords adds specific keywords to the abstract. These keywords may not seem obviously important to humans, but they can increase $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$ if added into the adversarial abstract. This phenomenon aligns with findings in the adversarial examples literature, where researchers have found the output of neural networks to be brittle to the insertion of certain keywords or tokens into the input. Furthermore, we find that the similarities can increase even when the keywords are used under different meanings and with different parts of speech across abstracts. This allows InsertKeywords to insert technical keywords from $\mathcal{Q}_{r}^{\text{adv}}$ into the $a^{\text{adv}}_{p}$ without introducing unrelated technical concepts that may confuse non-colluding readers or even raise suspicion. For example, “transfer learning” is a technical term, but the word “transfer” is also commonly used in English. To make sure that the abstract remains coherent and grammatically correct after keyword insertions, InsertKeywords may also make edits to parts of the abstract around the inserted keywords (e.g. insert a phrase that contains the keyword).
We present an example of InsertKeywords in Figure 3. In this example, the paper in adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$ proposes mitigating label scarcity in transfer learning with data augmentation, in particular using gradient flow methods to the minimize maximum mean discrepancy loss on the feature-Gaussian manifold. We use the FindKeywords subroutine to propose some keywords from $\mathcal{Q}_{r}^{\text{adv}}$ to add to the abstract $a_{p}$. Note that FindKeywords will return repeated keywords if adding multiple instances of the same word can further increase the similarity.
In some cases, keywords in the reviewer’s archive $\mathcal{Q}_{r}^{\text{adv}}$ are directly related to the abstract $a_{p}$, so these keywords can be inserted using the same meanings they have in $\mathcal{Q}_{r}^{\text{adv}}$. For the two instances of the ‘learning’ keyword, we insert the keyword instances by adding the phrase ‘machine learning’ twice, which match how the keyword ‘machine’ is used in the reviewer’s archive. Next, we add the keyword ‘feature-label’ into the abstract by breaking it up into a phrase ‘labels and features’. We also added the keyword ‘feature-Gaussian’ into the abstract but omitting the ‘feature-’ prefix and only inserting the word ‘Gaussian’. On the other hand, some keywords have specific technical meanings in the reviewer’s archive that are unrelated to this paper, but the keywords can still be inserted using their common English meanings. For the keyword ‘manifold’, we use it with the common English definition — a great deal, when the meaning used in the reviewer archive is feature-Gaussian ‘manifold’. We also use the keyword ‘discrepancy’ with its common English meaning — the quality or state of being different, when the meaning used in the reviewer’s archive is maximum mean ‘discrepancy’ loss. For the keyword ‘flow’, we insert it by adding the phrase ‘training flow’ referring to training pipelines, when the meaning used in the reviewer’s archive is gradient ‘flow’. However, keywords like ‘Riemannian’ and ‘optimum’ are considered as too technical and unrelated without common English meanings that would fit into the abstract $a_{p}$, so they would not be inserted.
@@
Example: With more real-world machine learning applications, the importance of safeguarding data privacy of labels and features has increased manifold. Per-example gradient clipping is a key algorithmic step that enables practical differential private (DP) training for deep learning models. The choice of clipping threshold $R$, however, is vital for avoiding high training loss discrepancy and achieving high accuracy under DP. Not only does it serve as a clipping threshold, but the Gaussian noises added are also dependent on $R$. We propose an easy-to-use replacement, called automatic clipping, that eliminates the need to tune $R$ for any DP optimizers, including DP-SGD, DP-Adam, DP-LAMB and many others. The automatic variants are as private and computationally efficient as existing DP optimizers, but require no DP-specific hyperparameters and thus improve DP training manifold. Our proposed method is as amenable as the standard non-private training flow. We give a rigorous convergence analysis of automatic DP-SGD in the non-convex setting, showing that it can enjoy an asymptotic convergence rate that matches the standard SGD, under a symmetric gradient noise assumption of the per-sample gradient (commonly used in the non-DP literature). We demonstrate on various language and vision tasks that automatic clipping outperforms or matches the state-of-the-art, and can be easily employed with minimal changes to existing machine learning codebases.
@@
Figure 3: An example of InsertKeywords modifying the abstract. In this example, the paper in adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$ proposes mitigating label scarcity in transfer learning with data augmentation, in particular using gradient flow methods to the minimize maximum mean discrepancy loss on the feature-Gaussian manifold. While some added keywords like ‘labels’, ‘features’, ‘learning’ and ‘Gaussian’ match their meanings used in reviewer’s archive $\mathcal{Q}_{r}^{\text{adv}}$, other keywords like ‘manifold’, ‘flow’, and ‘discrepancy’ take on their common English meanings. The InsertKeywords changes shown in this example, alongside adversarial reviewer archive curation, increase the reviewer’s similarity to the paper from being 101st most-similar to 3rd most-similar amongst all reviewers at the NeurIPS 2023 conference.
@@
Now, we describe the implementation details of the InsertKeywords operation. The process iteratively searches for ${\bm{M}}$ batches of keywords from the reviewer’s adversarial archive $\mathcal{Q}_{r}^{\text{adv}}$, inserting each batch of ${\bm{K}}$ keywords into $a^{\text{adv}}_{p}$ before searching for the next batch. Each batch of keywords is greedily chosen to help increase the similarity $s((t_{p},a^{\text{adv}}_{p}),\mathcal{Q}_{r}^{\text{adv}})$, as detailed by the FindKeywords subroutine (Algorithm 3) in Appendix A.2. Since FindKeywords is a greedy algorithm, the alternation between finding and inserting each batch of keywords can take into account the new $a^{\text{adv}}_{p}$ when finding the next batch of keywords. The two hyperparameters, ${\bm{M}}$ and ${\bm{K}}$, in InsertKeywords are tuned and reported in Section 5.1. The formalized algorithm of the InsertKeywords operation is detailed in Algorithm 2 in Appendix A.2.
@@
InsertKeywords also has human-in-the-loop and fully automatic modes (detailed in Algorithm 2). The differences between the two modes lie in how the keywords are inserted. In the human-in-the-loop mode, the human adds the keywords into $a^{\text{adv}}_{p}$ one by one. In addition, the human creates up to five drafts of different ways to add each keyword and keeps the draft with the maximum similarity. Since manually enforcing coherence and consistency constraints greatly limits where each keyword can be inserted in $a^{\text{adv}}_{p}$, we find that taking the maximum of multiple drafts is helpful, especially when the similarity is sensitive to the position each keyword is inserted at. On the other hand, the fully automatic mode asks a LLM to insert all keywords in the same batch at once. Again, the LLM is prompted to follow coherence and consistency constraints, but there are no human supervision to ensure them.

### v2

@@ 4 Attack Procedure
For any colluding paper-reviewer pair $(p,r)$, the attack’s goal is to have $r$ be assigned to $p$ for peer review at a conference. We begin by detailing the reviewer’s adversarial action in Section 4.1; then, we detail the author’s adversarial action in Section 4.2. Figure 1 is an overview of the procedure.
@@
(b) The colluding author adversarially modifies their abstract to be more similar to the paper in the colluding reviewer’s curated archive.
Figure 1: An illustrated example of the attack procedure. 1(a) shows the reviewer’s action; 1(b) shows the author’s action.
@@ 4.1 Adversarial Archive Curation by the Colluding Reviewer
Peer review assignment systems often allow reviewers to curate their own archives for text-matching. Some conferences even encourage this practice, starting with an empty archive that reviewers can then populate with any papers they choose. Platforms like OpenReview and the Toronto Paper Matching System (TPMS) [CZ13b] allow reviewers to manually add arbitrary publications by specifying titles and abstracts, or remove articles from their current profiles. While the ability to manually add arbitrary publications can amplify the effectiveness of attacks, our analysis focuses on the more common case where reviewers select papers from their existing profiles.
A colluding reviewer can exploit this process by constructing an adversarial archive, $Q_{r}^{\text{adv}}$, that includes only papers from their archive $Q_{r}$ which are highly similar to the target paper $p$. With max pooling, a commonly used aggregation strategy (Section 2.2), the reviewer only needs to ensure that their archive retains the paper most similar to $p$. Since max pooling considers only the highest similarity score, the presence of one highly similar paper is sufficient to bias the assignment, making this method particularly vulnerable.
@@
$Q_{r}^{\text{adv}}=\left\{\text{a random paper in }\operatorname*{argmax}_{q\in Q_{r}}\frac{\mathbf{v}_{p}\cdot\mathbf{v}_{q}}{\left\lVert\mathbf{v}_{p}\right\rVert_{2}\left\lVert\mathbf{v}_{q}\right\rVert_{2}}\right\},$ (1)
@@
For simplicity, the majority of our analyses focuses on the setting where the colluding reviewer has a single paper in their archive. Under max pooling (with arbitrary length archives), the similarity score obtained by the colluding reviewer for the target paper is greater than or equal to the similarity score obtained under our single paper setting. Thus, while these analyses of a single paper are not identical to max pooling, they give indications of attack potential under both max pooling and under reviewer-driven sub-selection under average pooling. We rigorously compare the two settings in Section 5.5 and find that max pooling is indeed more vulnerable.
@@ 4.2 Adversarial Abstract Modifications by the Colluding Author
Recall that only titles and abstracts are used to compute similarity scores between papers, and reviewers can curate their archives adversarially. In this section, we will expand notation for paper $p$ to have corresponding $t_{p}$ for the title and $a_{p}$ for the abstract as we will manipulate those, and we will explicitly denote the reviewer archive. When the archive has only one paper, max and average pooling are the same and we drop the subscript for brevity in notation. Hence, we use $s((t_{p},a_{p}),Q_{r})$ to denote paper-reviewer similarity below.
@@
To have a paper $p$ assigned to the colluding reviewer $r$ for peer review, the key idea behind the authors’ steps is to construct an adversarial abstract $a^{\text{adv}}_{p}$ to increase the similarity $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$ from $s((t_{p},a_{p}),Q_{r}^{\text{adv}})$ while maintaining consistency and coherence to avoid arousing suspicion.
We discover two simple abstract modification operations that effectively increase the similarity. The first is IncludeThemes which involves adding background or filler sentences related to the main ideas of papers in $Q_{r}^{\text{adv}}$. The second is InsertKeywords which inserts “keywords” that target SPECTER similarity to increase $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$ even if the keywords do not necessarily seem important to humans. This operation is inspired by works in adversarial robustness that show how small unexpected and unintelligible changes can break machine learning classifiers [ERLD17, JJZS20, LMG+20].
For each operation, we provide a description of the operation and an example modified abstract. Both operations employ Large Language Models (LLMs) and have two modes — human-in-the-loop and fully automatic. The human-in-the-loop mode is the way we expect colluding authors would modify their abstracts in reality, as human oversight can ensure there are no suspicious artifacts in the adversarial abstract. For several experiments in this paper, we use the fully automatic mode for scalability, but we also simulate and evaluate human-in-the-loop attacks at a feasible scale. The LLM model we use in this paper is the OpenAI gpt-4-0125-preview model.
Since adversarial abstract modification is a multi-step process, we also propose an optional early stopping check to stop further modifications if the attack is already envisaged to be successful. In one experiment, we use a realistic early stopping heuristic derived from past conference’s data: if the colluding reviewer is the most-similar reviewer for the colluding paper amongst all NeurIPS 2022 reviewers, no further abstract modifications are made. See Appendix A for details.
@@
The goal of IncludeThemes is to modify the abstract $a_{p}$ in order to increase the SPECTER similarity between paper $p$ and reviewer $r$, by adding background or filler sentences related to the central themes of papers in $Q_{r}^{\text{adv}}$. We note that the resulting modified abstract $a^{\text{adv}}_{p}$ may become inconsistent with its paper $p$ if the main ideas in the abstract have been changed entirely. Our key observation is that SPECTER similarities increase when abstracts share overlapping themes, even when those themes are central to papers in $Q_{r}^{\text{adv}}$ but only referenced as background information in $a^{\text{adv}}_{p}$. This means that IncludeThemes can produce $a^{\text{adv}}_{p}{}$ by including themes from $Q_{r}^{\text{adv}}$ to increase $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$ without violating the coherence and consistency constraints. An example of IncludeThemes can be found in Figure 2.
@@
Figure 2: An example of IncludeThemes modifying the abstract. Goal navigation is an important theme in the adversarial archive $Q_{r}^{\text{adv}}$ of the reviewer $r$. The modification (highlighted sentence) adds robot goal navigation as a motivating example for further improvements to multimodal models. The resulting abstract $a^{\text{adv}}_{p}$ remains coherent and consistent with the paper’s focus on text and image embedding alignment, but has increased similarity to the target adversarial archive $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$. The IncludeThemes changes shown in this example, alongside adversarial reviewer archive curation, increase the reviewer’s similarity to the paper from being 101st most-similar to become the most similar amongst all reviewers at the NeurIPS 2023 conference.
The implementation of the human-in-the-loop and fully automatic modes differ in whether human supervision is used to ensure the coherence and consistency of the modified abstract $a^{\text{adv}}_{p}$. In the human-in-the-loop mode, modified abstracts $a^{\text{adv}}_{p}$ are created by a human (potentially with the help of a LLM), and the human can make incremental edits to $a^{\text{adv}}_{p}$ to ensure coherence and consistency. On the other hand, the fully automatic mode only uses a LLM to generate the $a^{\text{adv}}_{p}$ with a prompt that emphasizes coherence and consistency, and it does not allow further edits to $a^{\text{adv}}_{p}$ once they are generated.
In addition, the implementation for both human-in-the-loop and fully automatic modes involves generating $\bm{N}$ different versions of modified abstracts. Since generating and modifying abstracts is stochastic (due to the stochasticity of LLM outputs and manual edits), we keep only the attempt that is the most similar to the colluding reviewer $r$. We tune and report the choice of $\bm{N}$ for our experiments in Section 5.1, and the formal algorithm is Algorithm 1 in Appendix A.1.
@@
Example: With more real-world machine learning applications, the importance of safeguarding data privacy of labels and features has increased manifold. Per-example gradient clipping is a key algorithmic step that enables practical differential private (DP) training for deep learning models. The choice of clipping threshold $R$, however, is vital for avoiding high training loss discrepancy and achieving high accuracy under DP. Not only does it serve as a clipping threshold, but the Gaussian noises added are also dependent on $R$. We propose an easy-to-use replacement, called automatic clipping, that eliminates the need to tune $R$ for any DP optimizers, including DP-SGD, DP-Adam, DP-LAMB and many others. The automatic variants are as private and computationally efficient as existing DP optimizers, but require no DP-specific hyperparameters and thus improve DP training manifold. Our proposed method is as amenable as the standard non-private training flow
@@
Figure 3: An example of InsertKeywords modifications. The paper in the adversarial archive $Q_{r}^{\text{adv}}$ proposes mitigating label scarcity in transfer learning with data augmentation. Some inserted keywords (highlighted) match their original meanings from the paper in $Q_{r}^{\text{adv}}$, while others (highlighted + bolded) are adapted to common English. The InsertKeywords changes to this abstract, alongside adversarial reviewer archive curation, increase the reviewer’s similarity to the paper from being 101st most-similar to 3rd most-similar amongst all reviewers at the NeurIPS 2023 conference.
Next, InsertKeywords adds specific keywords to the abstract. These keywords may not seem obviously important to humans, but they can increase $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$ if added into the adversarial abstract. This phenomenon aligns with findings in the adversarial examples literature; the outputs of neural networks can be brittle to the insertion of certain keywords or tokens into the input. Furthermore, we find that the similarities can increase even when the keywords are used under different meanings and with different parts of speech across abstracts. This allows InsertKeywords to insert technical keywords from $Q_{r}^{\text{adv}}$ into the $a^{\text{adv}}_{p}$ without introducing unrelated or suspicious concepts. For example, “transfer learning” might be simplified to “transfer,” a common English term. To ensure coherence and grammatical correctness, InsertKeywords may also adjust text around the inserted keywords (e.g. insert a phrase that contains the keyword).
We present an example of InsertKeywords in Figure 3. In this example, the paper in adversarial archive $Q_{r}^{\text{adv}}$ proposes mitigating label scarcity in transfer learning with data augmentation, in particular using gradient flow methods to the minimize maximum mean discrepancy loss on the feature-Gaussian manifold. We also propose the FindKeywords subroutine, which is described later, to suggests keywords from $Q_{r}^{\text{adv}}$ to add to the abstract $a_{p}$, including repeated words if they can further increase similarity.
Some keywords from the reviewer’s archive $Q_{r}^{\text{adv}}$ carry meanings directly related to the abstract $a_{p}$. In Figure 3, the keyword ‘learning’ is added as ‘machine learning’; ‘feature-label’ is broken up and added as ‘labels and features’; ‘feature-Gaussian’ is inserted only as ‘Gaussian’. For these keywords, the inserted variations match their meanings in the reviewer’s archive. However, other keywords with technical meanings unrelated to this paper $p$, such as ‘manifold’, ‘discrepancy’, and ‘flow’, are adapted to common English (‘manifold’ as “a great deal”, ‘discrepancy’ as “difference”, and ‘flow’ referring to training pipelines). Still, there can be unrelated technical keywords like ‘Riemannian’ and ‘optimum’ that do not have suitable common English meanings for abstract $a_{p}$, so they would not be inserted.
@@
Now, we describe the implementation details of the InsertKeywords operation. The process iteratively searches for $\bm{M}$} batches of $\bm{K}$} keywords, inserting each batch before searching for the next. We propose the FindKeywords subroutine (Algorithm 3 in Appendix A.2) to greedily choose keywords that increase the similarity $s((t_{p},a^{\text{adv}}_{p}),Q_{r}^{\text{adv}})$. Since FindKeywords is a greedy algorithm, the alternation between finding and inserting each batch of keywords can take into account the new $a^{\text{adv}}_{p}$ when finding the next batch of keywords. The two hyperparameters, $\bm{M}$} and $\bm{K}$}, in InsertKeywords are tuned and reported in Section 5.1. The formal algorithm of InsertKeywords operation is detailed in Algorithm 2 in Appendix A.2.
@@
InsertKeywords also has human-in-the-loop and fully automatic modes (detailed in Algorithm 2). The differences between the two modes lie in how the keywords are inserted. In the human-in-the-loop mode, the human adds the keywords into $a^{\text{adv}}_{p}$ one by one. In addition, the human creates up to five drafts of different ways to add each keyword and keeps the draft with the maximum similarity. Since manually enforcing coherence and consistency constraints greatly limits where each keyword can be inserted in $a^{\text{adv}}_{p}$, we find that taking the maximum of multiple drafts is helpful, especially when the similarity is sensitive to the position each keyword is inserted at. In the fully automatic mode, LLM inserts a whole batch of keywords each time, and the LLM is prompted to follow coherence and consistency constraints, but there are no human supervision to ensure them.